OpenStudy (anonymous):

Find the slope of the line that contains (5, -7) and (6, -4).

OpenStudy (anonymous):

-4 + 7 = 3 6 - 5 = 1 m = 3

OpenStudy (anonymous):

Y2-y1/x2-x1

OpenStudy (anonymous):

(x1,y1) -> (x2,y2) m = (y2 - y1) / (x2 - x1) (5,-7) (6,-4) -> your points plug them in as such m = (-4 - (-7)) / (6 - 5) m = 3/1 m = 3

OpenStudy (anonymous):

m = (-4 +7)/(6-5) = 3